OpenStudy (anonymous):

Subtracting a radical? Please help me!!!!! (7- √54) - (5 + √24)

OpenStudy (ajprincess):

\(\sqrt{54}=\sqrt{3*3*6}\) \(=3\sqrt{6}\) \(\sqrt{24}=\sqrt{2*2*6}\) \(=2\sqrt{6}\) \(7- \sqrt{54} - (5 + \sqrt{24})\) \(=7-3\sqrt{6}-5-2\sqrt{6}\) \(=7-5-3\sqrt{6}-2\sqrt{6}\) Can u do it @InsanelyChaotic?

OpenStudy (anonymous):

I'm attempting

OpenStudy (anonymous):

The answer would be \[2 - \sqrt{6}\]

OpenStudy (ajprincess):

First part of ur answrer is right \(-3\sqrt6-2\sqrt6=-5\sqrt6\)

OpenStudy (ajprincess):

Getting it @InsanelyChaotic?

OpenStudy (anonymous):

OH!!!! that's what I got the alst itme

OpenStudy (anonymous):

last*

OpenStudy (anonymous):

but i was told i was wrong....

OpenStudy (ajprincess):

guess u missed the 5. infront of \(\sqrt6\). right?

OpenStudy (ajprincess):

If ur answer was \(2-5\sqrt6\) then it is perfect:)

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

I was correct:) thank you!!!! :)))

OpenStudy (ajprincess):

welcome:)